Circular No. 8726

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 M3 (BARNARD) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation tabulated below) has been found to be cometary following posting on the ‘NEO Confirmation Page’. L. Buzzi (Varese, Italy; 0.60-m reflector) reports that images taken in good seeing during June 23.97–23.99 UT showed a 6′′ circular coma with a strong central condensation. Images from R. Salvo (Los Molinos, Uruguay) on June 24.2 and from J. Broughton (Reedy Creek, Queensland) on June 24.6 also suggest that the object is a comet. From a preliminary orbit computation, the undersigned suggested that the comet is a return of P/1889 M1 = 1889 III = 1889c (Barnard), for which A. Berberich (1889, A.N. 123, 77) was the first to compute an orbit of intermediate period (∼ 128 years). The orbital elements below are by B. G. Marsden, from observations 1889–2006 (mean residual 0′′.9), with astrometry also given on MPECs 2006-M38 and 2006-M42.

2006 UT  α_{2000}  δ_{2000}  Mag.
June 23.25870  17^{h}43^{m}38.^{s}68  −27^{o}26′53′′.5  17.1

Epoch = 1889 June 7.0 TT

T = 1889 June 21.2418 TT  ω = 60.^{o}1479
e = 0.953988  Ω = 272.5450 }2000.0
q = 1.102918 AU  i = 31.2189

a = 23.970332 AU n^{o} = 0.0083983 P = 117.36 years Epoch = 2006 Aug. 13.0 TT

T = 2006 Aug. 28.6903 TT  ω = 60.^{o}4661
e = 0.954400  Ω = 272.0666 }2000.0
q = 1.107142 AU  i = 31.2161

a = 24.279424 AU n^{o} = 0.0082385 P = 119.64 years

SUPERNOVA 2006di IN NGC 439 Further to IAUC 8723, L. A. G. Monard reports the discovery of an

apparent supernova (mag ∼ 16.1) on unfiltered CCD images taken on June 24.154 UT. The new object is located at α = 1^{h}13^{m}46^{s}.85, δ = −31^{o}44′30′′.2 (equinox 2000.0), which is 6′′ west and 19′′ north of the nucleus of NGC 439. Nothing is visible at this location on the Digitized Sky Survey (limiting red mag 20.5). Additional magnitudes from Monard for 2006di: June 1.137, [18.5; 25.134, 16.2; 26.137, 16.1.

2006 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green